# Sign changes of Kloosterman sums with moduli having at most two prime factors

### v1

@@ 3. Some Lemmas
4. Lower bound for $R_{1}(X)$
For the convenience of calculation, we only consider the sum restricted to the set as we mentioned in Lemma 3.1 which is
$P_{2}(X):=\left\{p_{1}p_{2}\sim X|X^{\eta}<p_{1}<p_{2},\ p_{1}>p_{2}^{3/4}X^{\eta}\right\}.$
This set corresponds to $R_{2}$ in the lemma 3.1 and $\eta$ is the same as in Lemma 3.1. Recall the definition of $\lambda_{d}$, in this case we see
$p_{1},p_{2}>\sqrt{D}.$
Therefore, a lower bound of $R_{1}(X)$ can be obatined, which is
$R_{1}(X)=$ $\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$\geq$ $\sum_{n\in P_{2}(X)}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\left(\sum_{d\mid\Pi_{l}}\lambda_{d}\right)^{2}\sum_{n\in P_{2}(X)}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|$
$\geq$ $(1+o(1))\left(\sum_{d\mid\Pi_{l}}\lambda_{d}\right)^{2}\times 4\tilde{g}(1)A_{2}(F)C_{2}\cdot\frac{X}{\log X}.$
In view of the selection of (1), we find that all $1/(\log X)^{j}$ vanishes in
$\left(\sum_{d\mid\Pi_{l}}\lambda_{d}\right)^{2},$
unless $j\geq 10$. For example, the term related to $1/\log X$ is equal to
$\frac{\binom{14}{1}}{\log\sqrt{D}}$ $\left(\log 2+\log 3+\dots+\log 29\right.$
$\left.-\log(2\cdot 3)-\log(2\cdot 29)-\dots-\log(23\cdot 29)-\dots-\log(2\cdot 3\cdot\dots\cdot 29)\right)=0.$
For the terms related to $1/(\log X)^{10}$, only
$10!\binom{14}{10}\left(\prod_{p\mid\Pi_{10}}\log p\right)\cdot\frac{1}{(\log\sqrt{D})^{10}}$
remains. Thus, we obtain the lower bound is
$(1+o(1))\left(\frac{10!\binom{14}{10}\prod_{p\mid\Pi_{10}}\log p}{(\log\sqrt{D})^{10}}\right)^{2}\times 4\tilde{g}(1)A_{2}(F)C_{2}\cdot\frac{X}{\log X}$
$=$ $(1+o(1))\left(4^{10}10!\binom{14}{10}\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 0.0142\tilde{g}(1)\frac{X}{(\log X)^{21}},$
where the numerical value of $A_{2}(F)=0.0319586...$ can be found in [[17], Proposition 2.1] or [[18], Section 4].
5. Upper bound for $R_{2}(X)$
We write $R_{2}(X)$ as
$R_{2}(X)=$ $\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\sum_{d\mid n\Pi_{l}}\xi_{d}$
$=$ $\sum_{d\leq D}\xi_{d}\sum_{n\Pi_{l}\equiv 0(d)}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n),$
@@ 5. Upper bound for $R_{2}(X)$
$\xi_{d}=\sum_{d=[h_{1},h_{2}]}\lambda_{h_{1}}\lambda_{h_{2}},\ |\xi_{d}|\leq 3^{\omega(d)}.$
The definition of $\xi_{d}$ indicates that $d$ is a square-free number. Let $d=d_{1}d_{2}$ satisfying $(d,\Pi_{l})=d_{1}$, we also have $(d_{2},\Pi_{l}/d_{1})=1$. Then $R_{2}(X)$ is
@@
$\sum_{d_{1}\mid\Pi_{l}}\sum_{d_{2}\leq D/d_{1}\atop(d_{2},\Pi_{l}/d_{1})=1}\xi_{d_{1}d_{2}}\sum_{n\equiv 0(d_{2})}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)$
@@
$\ll$ $\sum_{d_{1}\mid\Pi_{l}}3^{\omega(d_{1})}\sum_{d_{2}}3^{\omega(d_{2})}\left|\sum_{n\equiv 0(d_{2})}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\right|$
$\ll$ $\frac{X}{(\log X)^{A}}=o\left(\frac{X}{(\log X)^{21}}\right),$
where we adopt the Lemma 3.2. Thus Proposition 2.1 is proved.
6. Upper bound for $R_{3}(X)$

### v2

@@ 3. Some Lemmas
4. Proof of the second Proposition
Recalling the definition of $\tau(n;\alpha,\beta)$, we have
$\tau(n;\alpha,\beta)=\sum_{n=dl,d\in\mathcal{B}(n)\atop d\leq\sqrt{n}\text{exp}(-\sqrt{\log X})}\alpha^{\omega(d)}\beta^{\omega(l)}+\sum_{n=dl,d\in\mathcal{B}(n)\atop\sqrt{n}\text{exp}(-\sqrt{\log X})<d\leq\sqrt{n}}\alpha^{\omega(d)}\beta^{\omega(l)}.$
Then, from the twisted multiplicativity of Kloosterman sums and Weil’s bound, it’s beneficial to write
$R_{2}(X)=R_{21}(X)+R_{22}(X),$
corresponding to the two classes of $\tau(n;\alpha,\beta)$ mentioned above.
Firstly, we deal with $R_{22}(X)$. Setting
$\tau^{\prime}(n;\alpha,\beta):=\sum_{d\mid n,d\in\mathcal{B}(n)\atop\sqrt{n}\text{exp}(-\sqrt{\log X})<d\leq\sqrt{n}}\alpha^{\omega(d)}\beta^{\omega(l)},$
one can check that
$\tau^{\prime}(n;\alpha,\beta)\leq\sum_{d\mid n\atop\sqrt{n}\text{exp}(-\sqrt{\log X})<d\leq\sqrt{n}}\alpha^{\omega(d)}\beta^{\omega(l)}.$
We apply Weil’s bound of Kloosterman sums in $R_{22}(X)$. Therefore, our goal has become to deal with
(6) $R_{22}(X)\leq\sum_{\left(n,\Pi_{\varepsilon}\right)=1}g\left(\frac{n}{X}\right)\mu^{2}(n)\sum_{n=dl\atop\sqrt{n}\text{exp}(-\sqrt{\log X})<d\leq\sqrt{n}}\alpha^{\omega(2d)}\beta^{\omega(2l)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
The upper bound in (6) is the same as $R_{2}^{\sharp}(X)$ in Section 3 of reference [19], so we can directly obtain
$R_{22}(X)\leq X(\log X)^{-\frac{5}{4}}=o\left(\frac{X}{\log X}\right).$
Next we deal with $R_{31}(X)$. For $d\in\mathcal{B}(n)$, we know
$R_{21}(X)\leq\mathop{\sum\sum}_{m<n\atop{\omega(m)=3,\omega(n)\geq 3\atop\left(m,\Pi_{\varepsilon}\right)=1}}g\left(\frac{mn}{X}\right)\left|\text{Kl}(\overline{n}^{2};m)\right|\mu^{2}(mn)\alpha^{\omega(m)}(2\beta)^{\omega(n)}\left(\sum_{d\mid mn}\lambda_{d}\right)^{2},$
where we used Weil’s bound of Kloosterman sums and released the condition $\left(n,\Pi_{\varepsilon}\right)=1$.
Now, we follow the process in Section 4 of [17]. Firstly, we separate variables into small intervals by introducing a series of smooth functions. Setting
$\xi=1+(\log X)^{-B}$
for some parameter $B\geq 1$. Let $M$, NN be some parameters of the form
$M=\xi^{l_{1}},\ N=\xi^{l_{2}}$
for some $l_{1},l_{2}\in\mathbb{N}$. Thus, there exists sequences $\left\{b_{l_{1},\xi}\right\}_{l_{1}\geq 0}$, $\left\{b_{l_{2},\xi}\right\}_{l_{2}\geq 0}$ of smooth functions supported in $\xi^{l_{1}-1},\xi^{l_{1}+1}]$, $\xi^{l_{2}-1},\xi^{l_{2}+1}]$, respectively. And if we put
$U(x)=U_{l_{1}}(x)=b_{l_{1},\xi}(x),\ V(x)=V_{l_{2}}(x)=b_{l_{2},\xi}(x),$
then $U$ and $V$ satisfy
$\sum_{l_{1}\geq 0}U_{l_{1}}(x)=1,x\geq 1,$
$\sum_{l_{2}\geq 0}V_{l_{2}}(x)=1,x\geq 1.$
Also the derivatives of $U$ and $V$ satisfy
$x^{l}U^{(l)}(x),x^{l}V^{(l)}(x)\ll(\log X)^{Bl}$
for all $l\geq 0$. These can be guaranteed by Lemma 3.2 in [17].
Now, for $\textbf{H}=(M,N)$, $MN\sim X$, and $M\ll N$, we write
$R_{21}(X)\leq\sum_{\textbf{H}}R_{21}(X,\textbf{H}).$
Furthermore, we focus on the inner sum
$R_{21}(X,\textbf{H})=\sum_{l}\sum_{m\equiv 0(l)\atop\omega(m)=3,\left(m,\Pi_{\varepsilon}\right)=1}U(m)\mu^{2}(m)\alpha^{\omega(m)}\Psi(N;m,l),$
@@ 4. Proof of the second Proposition
$\Psi(N;m,l):=\sum_{(dn,m)=1\atop dn>m,\omega(dn)\geq 3}V(dn)g\left(\frac{mdn}{X}\right)\left|\text{Kl}(\overline{dn}^{2};m)\right|\mu^{2}(dn)(2\beta)^{\omega(dn)}\xi_{dl},$
with
$\xi_{n}:=\sum_{n=[d_{1},d_{2}]}\lambda_{d_{1}}\lambda_{d_{2}}.$
Set
$\Psi(N;m,l)=\Psi^{*}(N;m,l)-\Psi^{\sharp}(N;m,l),$
where
@@
$\Psi^{*}(N;m,l)=\sum_{(dn,m)=1\atop dn>m}V(dn)g\left(\frac{mdn}{X}\right)\left|\text{Kl}(\overline{dn}^{2};m)\right|\mu^{2}(dn)(2\beta)^{\omega(dn)}\xi_{dl},$
$\Psi^{\sharp}(N;m,l):=$ $\sum_{c_{1}+c_{2}<3}\sum_{(dn,m)=1\atop dn>m}V(dn)g\left(\frac{mdn}{X}\right)\left|\text{Kl}(\overline{dn}^{2};m)\right|\mu^{2}(dn)(2\beta)^{\omega(dn)}$
$\times\mathbbm{1}_{\omega(d)=c_{1}}\mathbbm{1}_{\omega(n)=c_{2}}\xi_{dl},$
with
$\mathbbm{1}_{\omega(q)=c}:=\begin{cases}&1,\ \text{if }\omega(q)=c,\\ &0,\ \text{if }\omega(q)\neq c.\end{cases}$
We then apply Lemma 3.3 and 3.4 to handle $R_{21}(X,\textbf{H})$. For further details, interested readers may refer to the treatment of $R_{2}^{\flat}(X)$ on Page 563 in [19].
For $\Psi^{*}(N;m,l)$, essentially, by invoking Lemma 3.3 with the assignments? $\boldsymbol{\alpha}=(2\beta)^{\omega(d)}$, $\boldsymbol{\beta}=(2\beta)^{\omega(n)}$, $\boldsymbol{\gamma}=\xi_{dl}$ ?and observing that the divisor function satisfies the “Siegel-Walfisz” condition, we derive the contribution of? $\Psi^{*}(N;m,l)$ to $R_{21}(X,\textbf{H})$ as
$\sum_{l}\sum_{m\equiv 0(l)\atop\omega(m)=3,\left(m,\Pi_{\varepsilon}\right)=1}U(m)\mu^{2}(m)\alpha^{\omega(m)}\Psi^{*}(N;m,l),$
where
$\Psi^{*}(N;m,l)=$ $\mathop{{\sum}^{*}}_{a(m)}\left|\text{Kl}(\overline{a}^{2};m)\right|\sum_{dn\equiv a(m)\atop dn>m}V(dn)g\left(\frac{mdn}{X}\right)\mu^{2}(dn)(2\beta)^{\omega(dn)}\xi_{dl}$
$=$ $(1+o(1))\frac{1}{\phi(m)}\mathop{{\sum}^{*}}_{a(m)}\left|\text{Kl}(\overline{a}^{2};m)\right|\sum_{(dn,m)=1\atop dn>m}V(dn)g\left(\frac{mdn}{X}\right)\mu^{2}(dn)(2\beta)^{\omega(dn)}\xi_{dl}.$
Recall that there are no small prime factors in the variable $m$. Now, we use Lemma 3.4 for the term $\mathop{{\sum}^{*}}_{a(m)}\left|\text{Kl}(\overline{a}^{2};m)\right|$ and obtain
$\mathop{{\sum}^{*}}_{a(m)}\left|\text{Kl}(\overline{a}^{2};m)\right|=\left(\frac{8}{3\pi}\right)^{\omega(m)}\phi(m)\prod_{p\mid m}(1+O(p^{-1/2}\log p))=(1+o(1))\left(\frac{8}{3\pi}\right)^{\omega(m)}\phi(m).$
The second equality holds because the prime factor $p$ is relatively large at this time, and their contribution is at most only $o\left(\frac{X}{\log X}\right)$. Therefore, we may derive $R_{21}(X,\textbf{H})$ with $\Psi^{*}(N;m,l)$ is bounded by
(7) $(1+o(1))\mathop{\sum\sum}_{m<n}U(m)V(n)g\left(\frac{mn}{X}\right)\mu^{2}(mn)\left(\frac{8\alpha}{3\pi}\right)^{\omega(m)}(2\beta)^{\omega(n)}\left(\sum_{d\mid mn}\lambda_{d}\right)^{2}.$
For $\Psi^{\sharp}(N;m,l)$, we see that there are at most two cases to deal with. Let
$\Psi^{\sharp}(N;m,l)=\Psi^{\sharp}_{1}(N;m,l)+\Psi^{\sharp}_{2}(N;m,l),$
where
$\Psi^{\sharp}_{1}(N;m,l):=$ $\sum_{c_{1}+c_{2}=1}\sum_{(dn,m)=1\atop dn>m}V(dn)g\left(\frac{mdn}{X}\right)\left|\text{Kl}(\overline{dn}^{2};m)\right|\mu^{2}(dn)(2\beta)^{\omega(dn)}$
$\times\mathbbm{1}_{\omega(d)=c_{1}}\mathbbm{1}_{\omega(n)=c_{2}}\xi_{dl},$
$\Psi^{\sharp}_{2}(N;m,l):=$ $\sum_{c_{1}+c_{2}=2}\sum_{(dn,m)=1\atop dn>m}V(dn)g\left(\frac{mdn}{X}\right)\left|\text{Kl}(\overline{dn}^{2};m)\right|\mu^{2}(dn)(2\beta)^{\omega(dn)}$
$\times\mathbbm{1}_{\omega(d)=c_{1}}\mathbbm{1}_{\omega(n)=c_{2}}\xi_{dl},$
respectively.
Thus, similar to the proof of $\Psi^{*}(N;m,l)$, we only need to replace the corresponding sequence in Lemma 3.3 with
$\boldsymbol{\alpha}=(2\beta)^{\omega(d)}\mathbbm{1}_{\omega(d)=1},\ \boldsymbol{\beta}=(2\beta)^{\omega(n)}\mathbbm{1}_{\omega(n)=1}\text{ or }(2\beta)^{\omega(n)}\mathbbm{1}_{\omega(n)=0},\ \boldsymbol{\gamma}=\xi_{dl}.$
At this point, $\boldsymbol{\alpha}$ is actually a prime sequence, so it still satisfies the “Siegel-Walfisz” condition through the application of Dirichlet’s prime theorem.
When applying Lemma 3.3, the condition $M\geq N$” requires our attention. In $\Psi^{*}(N;m,l)$ and $\Psi^{\sharp}(N;m,l)$, if the variable $d$ is less than $n$, we can directly apply Lemma 3.3. When the variable $d$ is larger than $n$, things will become slightly more complicated. In this case, we need to separate the variable $dl$ in $\xi_{dl}$ into independent variables to ensure that we can continue using Lemma 3.3.
According to the definition of $\xi_{dl}$, we have
$\xi_{dl}=$ $\sum_{dl=[d_{1},d_{2}]}\lambda_{d_{1}}\lambda_{d_{2}}$
$=$ $\sum_{dl=ke_{1}e_{2}}\lambda_{ke_{1}}\lambda_{ke_{2}}$
$=$ $\sum_{dl=ke_{1}e_{2}}\mu(ke_{1})\mu(ke_{2})G(\log k,\log e_{1},\log e_{2}),$
where $k=(d_{1},d_{2})$, $G$ is a polynomial (of $\log$) with the form
$\sum_{a,b_{1},b_{2}\geq 0}C_{a,b_{1},b_{2}}(\log k)^{a}(\log e_{1})^{b_{1}}(\log e_{2})^{b_{2}}.$
And $C_{a,b_{1},b_{2}}$ here is a constant coming from $F$ defined in Section 2. Recalling that all integers above are square-free numbers, thus we have decomposition
$d=k^{\prime}e_{1}^{\prime}e_{2}^{\prime},\ l=k^{\prime\prime}e_{1}^{\prime\prime}e_{2}^{\prime\prime}\ \text{by writing}\ k=k^{\prime}k^{\prime\prime},\ e_{1}=e_{1}^{\prime}e_{1}^{\prime\prime},\ e_{2}=e_{2}^{\prime}e_{2}^{\prime\prime}.$
Therefore, in $\Psi^{*}(N;m,l)$ or $\Psi^{\sharp}(N;m,l)$, we transform the sum into
$\sum_{d,n}\sum_{l\mid m}\text{(The summation term)}=\sum_{n,k^{\prime},k^{\prime\prime},e_{1}^{\prime},e_{1}^{\prime\prime},e_{2}^{\prime},e_{2}^{\prime\prime}}\ \ \sum_{(k^{\prime\prime}e_{1}^{\prime\prime}e_{2}^{\prime\prime})\mid m}\text{(The summation term)}.$
Consequently, $\xi_{dl}$ admits a multiplicative decomposition into factors depending solely on $d$ and $l$ respectively. At most, there will be an additional $(\log X)^{C}$ factor for some constant $C>0$, which comes from the polynomial $G$ and will contribute negligibly. This ensures that we can always use Lemma 3.3.
Combining with (7) and taking into account all admissible tuples H, we get
$R_{21}(X)\leq(1+o(1))\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\tau\left(n;\frac{8\alpha}{3\pi},2\beta\right)\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
From the definition of $\tau(n,X)$, we have
$\tau\left(n;\frac{8\alpha}{3\pi},2\beta\right)=$ $\left(\frac{8\alpha}{3\pi}\right)^{3}(2\beta)^{\omega(n)-3}N(n)$
$=$ $\left(\frac{4\alpha}{3\pi\beta}\right)^{3}(4\beta)^{\omega(n)}\frac{N(n)}{2^{\omega(n)}}$
(8) $\leq$ $\left(\frac{4\alpha}{3\pi\beta}\right)^{3}(4\beta)^{\omega(n)}\frac{\binom{\omega(n)}{3}}{2^{\omega(n)}},$
where
$N(n)=\#\left\{\mathcal{B}(n)\right\}.$
Recalling that
$\alpha>0,\ \beta>0,$
now we specifically take
(9) $\alpha=\alpha_{0}>0,\ \beta=1.$
The term
$\frac{\binom{\omega(n)}{3}}{2^{\omega(n)}}$
on the right side of (4) is decreasing for $\omega(n)\geq 6$, which leads to
$\tau\left(n;\frac{8\alpha}{3\pi},2\beta\right)\leq\frac{5}{16}\left(\frac{4\alpha_{0}}{3\pi}\right)^{3}4^{\omega(n)}.$
Then we have
$R_{2}(X)\leq\frac{5}{16}\left(\frac{4\alpha_{0}}{3\pi}\right)^{3}\sum_{\omega(n)\geq 6}g\left(\frac{n}{X}\right)\mu^{2}(n)4^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
Put
$F(x)=x^{4}(a_{0}+a_{1}x+a_{2}x^{2}+a_{3}x^{3}+a_{4}x^{4}).$
Similar with section 6 in [17], we shall minimize the proportion $c(k,F)/A_{2}(F)$. With the help of Mathematical software [^1] The code of the programs used can be found at https://github.com/MXZhong29/MXZhong.Github.io.git or requested from authors., we may choose
$(a_{0},a_{1},a_{2},a_{3},a_{4})\approx\left(\frac{100396}{53901},-\frac{17284}{13475},\frac{76486}{134753},-\frac{33241}{188654},\frac{10836}{377308}\right).$
Now we write
$\sum_{\omega(n)\geq 6}g\left(\frac{n}{X}\right)\mu^{2}(n)4^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}$
$=$ $\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)4^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}$
$-\sum_{i=1}^{5}\sum_{\omega(n)=i}g\left(\frac{n}{X}\right)\mu^{2}(n)4^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}$
$=$ $4\tilde{g}(1)c(4,F)\frac{X}{\log X}(1+o(1))$
$-\sum_{i=1}^{5}4^{i}\sum_{\omega(n)=i}g\left(\frac{n}{X}\right)\mu^{2}(n)\left(\sum_{d\mid n}\lambda_{d}\right)^{2},$
through the application of Lemma 3.5.
Furthermore, by using prime number theorem (or see [10], Page 293), we get
$\sum_{\omega(n)=i}g\left(\frac{n}{X}\right)\mu^{2}(n)\left(\sum_{d\mid n}\lambda_{d}\right)^{2}$
$=$ $(1+o(1))B_{i}(F)\widetilde{g}(1)\frac{X}{\log X},$
where $B_{1}(F)=1$, and for $i\geq 2$,
$B_{i}(F)=\mathop{\int\ldots\int}_{T_{i}}\frac{L_{i}^{2}\left(F;X^{1-\alpha_{2}-\cdots-\alpha_{i}},X^{\alpha_{2}},\ldots,X^{\alpha_{i}}\right)}{\alpha_{2}\cdots\alpha_{i}(1-\alpha_{2}-\cdots-\alpha_{i})}\text{d}\alpha_{2}\cdots\text{d}\alpha_{i},$
$L_{i}(F;\alpha_{1},\alpha_{2},\ldots,\alpha_{i})=\sum_{\mathcal{A}\subseteq\left\{\alpha_{1},\alpha_{2},\ldots,\alpha_{i}\right\}\atop\sum_{\alpha\in\mathcal{A}}\alpha<\frac{1}{4}}(-1)^{|\mathcal{A}|}F\left(1-4\sum_{\alpha\in\mathcal{A}}\alpha\right),$
with $T_{i}:=\left\{(\alpha_{2},\ldots,\alpha_{i})\in[10^{-2025},1)^{i-1}:\alpha_{i}<\cdots<\alpha_{2}<1-\alpha_{2}-\cdots-\alpha_{i}\right\}$. Then
$\sum_{\omega(n)\geq 6}g\left(\frac{n}{X}\right)\mu^{2}(n)4^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}$
$=$ $(1+o(1))\widetilde{g}(1)\left(4c(4,F)-\sum_{i=1}^{5}4^{i}B_{i}(F)\right)\frac{X}{\log X}.$
Recalling the value of $F(x)$, with the help of mathematical software, we obtain
$R_{2}(X)\leq(1+o(1))6.27044\alpha_{0}^{3}\times\widetilde{g}(1)\frac{X}{\log X}.$
5. Proof of the first Proposition
We divide $R_{1}(X)$ into two parts, namely
$R_{1}(X)=R_{11}(X)-R_{12}(X),$
where
$R_{11}(X)=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left(\left|\text{Kl}(1;n)\right|\pm\text{Kl}(1;n)\right)\left(\sum_{d\mid n}\lambda_{d}\right)^{2},$
$R_{12}(X)=\sum_{\left(n,\Pi_{\varepsilon}\right)>1}g\left(\frac{n}{X}\right)\mu^{2}(n)\left(\left|\text{Kl}(1;n)\right|\pm\text{Kl}(1;n)\right)\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
Lemma 3.2 tells us that
$R_{11}(X)=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n}\lambda_{d}\right)^{2}+o\left(\frac{X}{\log X}\right).$
Interested readers may see Section 6 in [16] for the details, where the Kloosterman sums’ spectral theory was used.
Recall the value of $F(x)$ again. Then from Lemma 3.1 we have
$R_{11}(X)\geq(1+o(1))0.76235\times\widetilde{g}(1)\frac{X}{\log X}.$
Now we turn to handle $R_{12}(X)$. One may notice that $R_{12}(X)$ here is the same as $R_{12}(X)$ in [19], but we will not directly cite this result. We use a similar approach to handle $R_{12}(X)$, which is more direct compared to the method in [19]. The approach we use here can avoid using Cauchy’s inequality and avoid introducing a new smooth function $V(X)$.
Naturally, we have
$R_{12}(X)\leq 2\sum_{\left(n,\Pi_{\varepsilon}\right)>1}g\left(\frac{n}{X}\right)\mu^{2}(n)2^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2},$
where we used the Weil’s bound for Kloosterman sums. Furthermore, we can write
$R_{12}(X)\leq$ $2\sum_{p<X^{\varepsilon}}\sum_{n\equiv 0(p)}g\left(\frac{n}{X}\right)\mu^{2}(n)2^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}$
(10) $=$ $4\sum_{p<X^{\varepsilon}}\sum_{(n,p)=1}g\left(\frac{np}{X}\right)\mu^{2}(n)2^{\omega(n)}\left(\sum_{d\mid np}\lambda_{d}\right)^{2}.$
Generally, for the inner sum, we will consider the following form of sum
$T(X):=\sum_{(n,\Pi_{l})=1}g\left(\frac{n\Pi_{l}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2},$
where we set
$\Pi_{l}=\prod_{1\leq i\leq l}p_{i}.$